# LEAF: Unveiling Two Sides of the Same Coin in Semi-supervised Facial Expression Recognition

### v1

@@ 5. Experiments
Pi Model (Laine and Aila, 2016) 43.39 52.92 60.29 42.62 48.18 58.06 44.76 50.34 53.26 39.71 44.26 46.88
Pseudo-Label (Lee et al., 2013) 44.36 52.04 59.61 39.75 48.03 55.19 44.85 48.28 52.28 40.70 43.03 47.49
VAT (Miyato et al., 2018) 33.28 48.21 57.14 40.07 48.40 51.95 43.46 47.68 51.79 38.02 43.13 47.54
@@
MeanTeacher (Tarvainen and Valpola, 2017) 39.50 52.26 62.71 37.86 47.43 58.40 46.98 50.47 54.09 41.76 45.87 49.57
MixMatch (Berthelot et al., 2019b) 40.40 54.77 62.62 43.72 51.06 56.64 45.20 49.12 53.82 41.12 44.07 49.05
ReMixMatch (Berthelot et al., 2019a) 39.52 54.22 60.93 40.63 48.30 58.15 43.63 50.80 54.02 42.66 45.34 50.17
FixMatch (Sohn et al., 2020) 48.13 60.23 65.04 48.05 52.77 60.38 47.83 51.39 54.75 42.69 46.04 49.50
DeFixMatch (Schmutz et al., 2022) 50.37 59.21 65.81 45.08 53.01 62.14 47.51 51.32 55.00 43.28 46.49 48.49
Dash (Xu et al., 2021) 49.95 59.27 66.44 43.10 55.56 62.74 48.03 51.65 54.49 42.33 45.69 50.06
CoMatch (Li et al., 2021b) 49.09 60.81 65.18 41.24 49.32 59.74 47.97 51.88 55.07 43.94 46.47 49.03
@@
AdaMatch (Berthelot et al., 2021) 48.35 57.10 63.61 43.87 50.48 59.92 47.03 51.13 54.23 41.69 45.43 49.12
FlexMatch (Zhang et al., 2021) 48.79 53.24 61.16 43.75 47.90 58.23 48.00 52.23 54.76 42.98 46.45 49.02
FreeMatch (Wang et al., 2022) 48.89 55.26 62.09 41.85 45.57 58.39 46.74 50.77 54.46 41.98 45.96 49.44
SoftMatch (Chen et al., 2023) 49.95 56.13 63.45 42.97 49.97 57.58 47.65 50.71 54.56 42.06 46.73 49.26
@@
Figure 4. Performance comparison about overall accuracy and balanced accuracy with respect to different numbers of labels.
5.1. Datasets and Evaluation Metrics
@@
We conduct extensive experiments on several widely used benchmark datasets that vary in granularity and size. RAFDB (Li et al., 2017) consists of facial images from seven distinct expressions. It is divided into a training set containing 12,271 images and a testing set with 3,068 images. FERPlus (Barsoum et al., 2016) comprises eight facial expressions, with a total of 24,941 training images, 3,589 validation images, and 3,589 test images. AffectNet (Mollahosseini et al., 2017) is a large-scale dataset containing 286,564 training images and 4,000 test images, all manually annotated with eight expression labels. We consider two versions of AffectNet: AffectNet7 and AffectNet8. AffectNet7 excludes the expression category of contempt, which consists of 3,667 training images and 500 test images.
@@
Considering the imbalanced nature of the datasets, we argue that balanced accuracy is a more appropriate evaluation metric for the FER task than overall accuracy. Balanced accuracy takes into account the successful recognition of both major and minor classes (Zhang et al., 2024), providing a more comprehensive assessment of the model’s performance. Unless otherwise specified, all reported results in this paper are based on balanced accuracy.
@@
Table 2. Ablation on the structure of experts across various scenarios. The best results are shown in boldface.
@@
Table 3. Ablation on proposed components on AffectNet7 with different settings. The best results are shown in boldface.
@@
5.2. Implementation Details
We implement our proposed LEAF using the PyTorch framework and conduct all experiments on NVIDIA Tesla A100 GPUs. The facial images are aligned and resized to $224\times 224$ pixels. We employ the Adam optimizer with a learning rate of $45e-$ 4, a batch size of 64, and train the models for 20 epochs. For a fair comparison, we select ResNet-18 (He et al., 2016) pre-trained on MS-Celeb-1M (Guo et al., 2016) as the backbone network for all baselines and our proposed method. The hyperparameters of each baseline are set according to their respective papers, and we extend these methods to the FER domain. It is important to note that we do not compare our LEAF with fully supervised FER methods, as these approaches require large amounts of labeled data and are unable to utilize unlabeled facial expression samples to improve recognition performance.
5.3. Quantitative Comparison
Tab. 1 presents a comprehensive quantitative comparison of our proposed LEAF with state-of-the-art semi-supervised learning methods on four widely used FER benchmark datasets. From these results, we can draw several key observations. Firstly, it is evident that semi-supervised FER methods, such as AdaCM (Li et al., 2022) and our LEAF, outperform semi-supervised image classification methods in most cases. This can be attributed to the fact that these FER-specific approaches address the unique challenges of subtle inter-class differences among facial expressions. By effectively utilizing unlabeled data through techniques like contrastive learning for low-confidence samples and hierarchical integration of EAF modules, these methods can extract more discriminative representations and improve overall performance. Secondly, we observe that the recognition of seven emotions generally yields better results compared to the recognition of eight emotions in the semi-supervised setting. This can be explained by the subtle distinction between the newly added emotion of contempt and the existing basic emotions of anger and disgust. Distinguishing between these emotions when limited labeled data is available poses a significant challenge, as the nuances can be difficult to capture without sufficient training examples. Finally, and most importantly, our proposed LEAF consistently outperforms all other approaches across various settings on the four benchmark datasets. This impressive performance can be attributed to several key components of our method. The semantic-level and instance-level EAF modules play a crucial role in extracting expression-relevant representations, while the category-level EAF strategy enables the assignment of propoer pseudo-labels to ambiguous samples. The synergistic combination of these modules allows LEAF to effectively leverage both labeled and unlabeled data, resulting in robust and superior performance. A more detailed analysis of each module is provided in Sec. 5.5.2.
5.4. Qualitative Analysis
@@
Figure 5. The t-SNE visualization with 1600 labeled samples on RAFDB and FERPlus.
Figure 6. Sensitivity analysis of the number of experts with 100 labeled samples on RAFDB and FERPlus.
5.5. Ablation Studies and Analysis
5.5.1. Impact of Expert Structure
@@
5.5.2. Contribution of Proposed Components
@@
5.5.3. Sensitivity to the Number of Experts
@@
Figure 7. Performance before and after integrating our EAF strategies into other methods on RAFDB.
5.5.4. Integration with Other Semi-Supervised Methods
To demonstrate the flexibility and effectiveness of our proposed EAF strategies, we integrate them with two state-of-the-art semi-supervised methods, FixMatch (Sohn et al., 2020) and FlexMatch (Zhang et al., 2021). The results, presented in Fig. 7, show that our EAF strategies consistently bring improvements when the label quantity varies from $100$ to $300$. Notably, when the number of labels is extremely small (i.e., 100 labels), the EAF strategies provide even greater improvements, highlighting their effectiveness in scenarios with severely limited labeled data.
5.6. Visualization Analysis
To gain further insights into the learned feature representations, we perform t-SNE (Van der Maaten and Hinton, 2008) visualization of the features extracted from the encoder. Fig. 5 presents the visualization results, where each facial expression is represented with a different color. The visualization results demonstrate that even under label scarcity, both AdaCM (Li et al., 2022) and our LEAF are able to maintain good discriminability between different expressions. This indicates that these methods are capable of learning meaningful and separable feature representations for facial expressions, despite the limited availability of labeled data. It is important to note that AdaCM (Li et al., 2022) achieves this discriminability by employing a contrastive learning objective, which aims to increase the distances between different expressions in the feature space. In contrast, our LEAF utilizes EAF strategies to focus the network’s attention on expression-related information. These two distinct representation learning strategies demonstrate the effectiveness of different approaches in tackling the FER task. Upon closer inspection of the visualization results, we observe that our approach slightly outperforms AdaCM (Li et al., 2022) in terms of the compactness and separability of the learned feature representations. This suggests that the hierarchical decoupling and fusing strategy employed by LEAF, along with the ambiguous pseudo-label generation strategy, enables the model to capture more discriminative and expression-relevant representations.
6. Conclusion

### v3

@@ 5 Experiments
Pi Model Laine and Aila (2016) 43.39 52.92 60.29 42.62 48.18 58.06 44.76 50.34 53.26 39.71 44.26 46.88
Pseudo-Label Lee et al. (2013) 44.36 52.04 59.61 39.75 48.03 55.19 44.85 48.28 52.28 40.70 43.03 47.49
VAT Miyato et al. (2018) 33.28 48.21 57.14 40.07 48.40 51.95 43.46 47.68 51.79 38.02 43.13 47.54
UDA Xie et al. (2020) 49.22 59.96 65.80 44.36 52.48 62.80 47.69 51.49 55.48 42.74 46.38 48.73
MeanTeacher Tarvainen and Valpola (2017) 39.50 52.26 62.71 37.86 47.43 58.40 46.98 50.47 54.09 41.76 45.87 49.57
MixMatch Berthelot et al. (2019) 40.40 54.77 62.62 43.72 51.06 56.64 45.20 49.12 53.82 41.12 44.07 49.05
ReMixMatch Berthelot et al. (2019) 39.52 54.22 60.93 40.63 48.30 58.15 43.63 50.80 54.02 42.66 45.34 50.17
FixMatch Sohn et al. (2020) 48.13 60.23 65.04 48.05 52.77 60.38 47.83 51.39 54.75 42.69 46.04 49.50
DeFixMatch Schmutz et al. (2022) 50.37 59.21 65.81 45.08 53.01 62.14 47.51 51.32 55.00 43.28 46.49 48.49
Dash Xu et al. (2021) 49.95 59.27 66.44 43.10 55.56 62.74 48.03 51.65 54.49 42.33 45.69 50.06
CoMatch Li et al. (2021) 49.09 60.81 65.18 41.24 49.32 59.74 47.97 51.88 55.07 43.94 46.47 49.03
@@
AdaMatch Berthelot et al. (2021) 48.35 57.10 63.61 43.87 50.48 59.92 47.03 51.13 54.23 41.69 45.43 49.12
FlexMatch Zhang et al. (2021) 48.79 53.24 61.16 43.75 47.90 58.23 48.00 52.23 54.76 42.98 46.45 49.02
FreeMatch Wang et al. (2022) 48.89 55.26 62.09 41.85 45.57 58.39 46.74 50.77 54.46 41.98 45.96 49.44
SoftMatch Chen et al. (2023) 49.95 56.13 63.45 42.97 49.97 57.58 47.65 50.71 54.56 42.06 46.73 49.26
@@
Figure 4: Performance comparison about overall accuracy and balanced accuracy with respect to different numbers of labels.
5.1 Datasets and Evaluation Metrics
@@
We conduct extensive experiments on several widely used benchmark datasets that vary in granularity and size. RAFDB Li et al. (2017) consists of facial images from seven distinct expressions. It is divided into a training set containing 12,271 images and a testing set with 3,068 images. FERPlus Barsoum et al. (2016) comprises eight facial expressions, with a total of 24,941 training images, 3,589 validation images, and 3,589 test images. AffectNet Mollahosseini et al. (2017) is a large-scale dataset containing 286,564 training images and 4,000 test images, all manually annotated with eight expression labels. We consider two versions of AffectNet: AffectNet7 and AffectNet8. AffectNet7 excludes the expression category of contempt, which consists of 3,667 training images and 500 test images.
@@
Considering the imbalanced nature of the datasets, we argue that balanced accuracy is a more appropriate evaluation metric for the FER task than overall accuracy. Balanced accuracy takes into account the successful recognition of both major and minor classes Zhang et al. (2024), providing a more comprehensive assessment of the model’s performance. Unless otherwise specified, all reported results in this paper are based on balanced accuracy.
5.2 Implementation Details
We implement our proposed LEAF using the PyTorch framework and conduct all experiments on NVIDIA Tesla A100 GPUs. The facial images are aligned and resized to $224\times 224$ pixels. We employ the Adam optimizer with a learning rate of $45e-$ 4, a batch size of 64, and train the models for 20 epochs. For a fair comparison, we select ResNet-18 He et al. (2016) pre-trained on MS-Celeb-1M Guo et al. (2016) as the backbone network for all baselines and our proposed method. The hyperparameters of each baseline are set according to their respective papers, and we extend these methods to the FER domain. It is important to note that we do not compare our LEAF with fully supervised FER methods, as these approaches require large amounts of labeled data and are unable to utilize unlabeled facial expression samples to improve recognition performance.
5.3 Quantitative Comparison
Tab. 1 presents a comprehensive quantitative comparison of our proposed LEAF with state-of-the-art semi-supervised learning methods on four widely used FER benchmark datasets. From these results, we can draw several key observations. Firstly, it is evident that semi-supervised FER methods, such as AdaCM Li et al. (2022), LION Du et al. (2023), and our LEAF, outperform semi-supervised image classification methods in most cases. This can be attributed to the fact that these FER-specific approaches address the unique challenges of subtle inter-class differences among facial expressions. By effectively utilizing unlabeled data through techniques like contrastive learning for low-confidence samples and hierarchical integration of EAF modules, these methods can extract more discriminative representations and improve overall performance. Secondly, we observe that the recognition of seven emotions generally yields better results compared to the recognition of eight emotions in the semi-supervised setting. This can be explained by the subtle distinction between the newly added emotion of contempt and the existing basic emotions of anger and disgust. Distinguishing between these emotions when limited labeled data is available poses a significant challenge, as the nuances can be difficult to capture without sufficient training examples. Finally, and most importantly, our proposed LEAF consistently outperforms all other approaches across various settings on the four benchmark datasets. This impressive performance can be attributed to several key components of our method. The semantic-level and instance-level EAF modules play a crucial role in extracting expression-relevant representations, while the category-level EAF strategy enables the assignment of proper pseudo-labels to ambiguous samples. The synergistic combination of these modules allows LEAF to effectively leverage both labeled and unlabeled data, resulting in robust and superior performance. A more detailed analysis of each module is provided in Sec. 5.5.2.
@@ 5.3 Quantitative Comparison
Table 2: Ablation on the structure of experts across various scenarios. The best results are shown in boldface.
@@
Table 3: Ablation on proposed components on AffectNet7 with different settings. The best results are shown in boldface.
@@
5.4 Qualitative Analysis
@@
Figure 5: The t-SNE visualization with 1600 labeled samples on RAFDB and FERPlus.
Figure 6: Sensitivity analysis of the number of experts with 100 labeled samples on RAFDB and FERPlus.
5.5 Ablation Studies and Analysis
5.5.1 Impact of Expert Structure
@@
5.5.2 Contribution of Proposed Components
@@
5.5.3 Sensitivity to the Number of Experts
@@
Figure 7: Performance before and after integrating our EAF strategies into other methods on RAFDB.
5.5.4 Integration with Other Semi-Supervised Methods
To demonstrate the flexibility and effectiveness of our proposed EAF strategies, we integrate them with two state-of-the-art semi-supervised methods, FixMatch Sohn et al. (2020) and FlexMatch Zhang et al. (2021). The results, presented in Fig. 7, show that our EAF strategies consistently bring improvements when the label quantity varies from $100$ to $300$. Notably, when the number of labels is extremely small (i.e., 100 labels), the EAF strategies provide even greater improvements, highlighting their effectiveness in scenarios with severely limited labeled data.
5.6 Visualization Analysis
To gain further insights into the learned feature representations, we perform t-SNE Van der Maaten and Hinton (2008) visualization of the features extracted from the encoder. Fig. 5 presents the visualization results, where each facial expression is represented with a different color. The visualization results demonstrate that even under label scarcity, both AdaCM Li et al. (2022) and our LEAF are able to maintain good discriminability between different expressions. This indicates that these methods are capable of learning meaningful and separable feature representations for facial expressions, despite the limited availability of labeled data. It is important to note that AdaCM Li et al. (2022) achieves this discriminability by employing a contrastive learning objective, which aims to increase the distances between different expressions in the feature space. In contrast, our LEAF utilizes EAF strategies to focus the network’s attention on expression-related information. These two distinct representation learning strategies demonstrate the effectiveness of different approaches in tackling the FER task. Upon closer inspection of the visualization results, we observe that our approach slightly outperforms AdaCM Li et al. (2022) in terms of the compactness and separability of the learned feature representations. This suggests that the hierarchical decoupling and fusing strategy employed by LEAF, along with the ambiguous pseudo-label generation strategy, enables the model to capture more discriminative and expression-relevant representations.
6 Conclusion